OpenStudy (anonymous):

Solve the following system of equations. Please show your work. x - y = 13 2x + y = 2

myininaya (myininaya):

Add the equations together You have (x+2x)+(-y+y)=(13+2) Solve this equation for x Then go back to one of the equations given and plug that x in and solve for y

OpenStudy (anonymous):

I don't understand

myininaya (myininaya):

Adding part?

jhonyy9 (jhonyy9):

so than will result 3x=15 x=15/3 x=5 5-y=13 y=-8 x=5 y=-8